# Self interacting scalar field theory in general curved spacetimes at zero and finite temperature revisited

### v1

@@ 5 Spontaneous symmetry breaking and phase transition at zero temperature
The chief motivation of this section is to see the effect of the spacetime curvature in the phase transition or spontaneous symmetry breaking phenomenon, at zero temperature. As we have already mentioned, this topic is old. In flat spacetime, one of the earliest thermal field theory calculations in this context can be seen in the seminal work [7]. At zero temperature and with spacetime curvature some such computations can be seen in e.g. [4, 33, 38] and references therein. In particular, [33, 38] use weak or linear in curvature approximation and renormalisation group techniques. However as we have also mentioned earlier, such results cannot be applicable to Ricci flat spacetimes.
@@
Let us now consider the zero temperature effective potential in the absence of curvature, 20 ($R=0=f_{1}$), in the dimensionless form
@@
$\left.+\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right]$ (40)
Putting in now some customary numerical values, $m\sim 0.00051{\rm GeV}$ and $\mu\sim 3\times 10^{-25}{\rm GeV}$, which is appropriate for the large scale physics we are interested in, e.g. [80] and references therein, we have plotted the above zero temperature flat space effective potential in 3. We note that since the scalar we have taken has non-vanishing rest mass, there can be no spontaneous symmetry breaking for the quartic self interaction. However, in the presence of spacetime curvature, the scenario can change, as we wish to show below.
@@
Figure 3: The one loop effective potential, 40, with vanishing temperature in flat spacetime. Note that for $\bar{\eta}=0$, no spontaneous symmetry breaking is indicated, as we have taken the rest mass of the scalar to be non-vanishing.
@@
in 20, so that we have around $r\sim 2GM$
@@
$\left.+\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right]+\frac{0.0016}{m^{4}G^{4}M^{4}}\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)$ (41)
Note that the effective potential diverges as $M\to 0$. This corresponds to the fact that the value of $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ increases on the horizon with decreasing mass of the black hole. On the other hand, the result approaches the flat spacetime limit, 40, for large value of the black hole mass. Taking now for example, $GM\sim 10^{2}{\rm m}$, we have plotted 41 in 4. We note in particular for the first of 4, that for increasing cubic coupling, the effective potential develops a deep. This originates from the logarithm contained in the last term on the right hand side of 41, which may contribute a large negative value. Note also that for a given $\lambda$ value, we cannot arbitrarily go on increasing the cubic coupling, owing to the discussion appearing at the end of the preceding Section, 39.
@@
Figure 4: The one loop effective potential, 41, at vanishing temperature in the Schwarzschild spacetime. See main text for discussion.
@@
$+\left(r^{2}+a^{2}+\frac{2a^{2}GMr\sin^{2}\theta}{r^{2}+a^{2}\cos^{2}\theta}\right)\sin^{2}\theta d\phi^{2}$ (42)
@@
$R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}=\frac{48G^{2}M^{2}\left(r^{6}-15a^{2}r^{4}\cos^{2}\theta+15a^{4}r^{2}\cos^{2}\theta-a^{6}\cos^{6}\theta\right)}{\left(r^{2}+a^{2}\cos^{2}\theta\right)^{6}}$
Comparing this with the Schwarzschild spacetime, we see that there is no qualitative change at all due to spacetime rotation on the equatorial plane, $\theta=\pi/2$. The effect of the spacetime rotation is maximum on the axis ($\theta=0,\pi$). We note that if we take the extremal limit ($a\to GM$), the Kretschmann scalar vanishes as we move towards the horizon ($r\to GM$) along the axis. This shows that the one loop effective potential for an extreme Kerr black hole spacetime in the vicinity of the axial points near the horizon coincides with that of the flat spacetime, 40.
@@
Let us now come to the de Sitter spacetime,
$ds^{2}=-dt^{2}+e^{2\sqrt{\Lambda}t/\sqrt{3}}\left(dx^{2}+dy^{2}+dz^{2}\right)$
for which we have $R=4\Lambda$, $R^{\mu\nu}R_{\mu\nu}=4\Lambda^{2}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=8\Lambda^{2}/3$ in 20, $\Lambda$ being the positive cosmological constant. We have the one loop effective potential
@@
$+\frac{0.032\Lambda^{2}}{m^{4}}\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\left[\left(\frac{1}{6}-\xi\right)^{2}-0.001\right]$ (43)
We have investigated various features of this potential in 5, 6 and 7.
@@
Figure 5: The one loop effective potential, 43, at vanishing temperature in the de Sitter spacetime. Comparing the latter with 3(b), 4(b), we see that spontaneous symmetry breaking in de Sitter with a massive scalar with $m^{2}{>}0$ is possible. In both the plots appearing above we have taken $\Lambda\sim 10^{-52}{\rm m}^{-2}$, i.e. its current observed value and $\xi\sim 0.13$. See main text for discussion.
In particular, we note that (with $\bar{\eta}=0$), spontaneous symmetry breaking at one loop is possible in de Sitter even with $m^{2}{>}0$. This has no analogue for flat or the Ricci flat spacetimes we have discussed, 3(b), 4(b). In fact symmetry breaking for a massive scalar in flat spacetime seems only to be possible if the scale $\mu$ is extremely large compared to $m$. In de Sitter, this effect originates from the term appearing in the last line of the right hand side of 43, which, when $\xi$ is $1/6$ or close to it, yields a negative term which can dominate the $\bar{\phi}^{2}/2$ term. Eventually this leads to the symmetry breaking effective potential. We see that this is possible with a $\Lambda$ as tiny as its current observed value. Certainly, such effect may a priori be expected to be much more stronger in the early inflationary universe scenario, where the dark energy density was much higher compared to that of today. Note also from the second of 7 that symmetry breaking remains possible for a very wide range of the renormalisation scale $\mu$. Finally, we note that had we ignored the quadratic curvature terms from the beginning, we would not have seen any such effect. We believe this feature, along with 27, sufficiently motivate us to look into this problem by employing some non-perturbative resummation techniques. Such technique should not only resum the coupling, but must resum some curvature terms as well. We reserve this problem for a future work.
@@
Figure 6: The one loop effective potential, 43, at vanishing temperature in the de Sitter spacetime for different values of the non-minimal coupling $\xi$. Note that for $\xi\sim 0.167$ or $1/6$, the dip or the hill of the effective potential is maximum. This effect originates from the terms appearing on the last line on the right hand side of 43. See main text for discussion.
@@
Figure 7: The one loop effective potential, 43, with vanishing temperature in the de Sitter spacetime for different values of the cosmological constant $\Lambda$, and the mass scale $\mu$. Increasing the value of $\Lambda$ increases the strength of the symmetry breaking effect. Note also that the symmetry breaking feature of the potential remains valid for a very wide range of the renormalisation scale, $\mu$.
@@ 5.1 Phase transition at finite temperature
To begin with, let us consider the thermal effective potential 35 at zero spacetime curvature with only quartic coupling and at high temperature, e.g. [27],
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}+\frac{\lambda\phi^{4}}{4!}+\left[-\frac{\pi^{2}}{90\beta^{4}}+\frac{m^{2}+\frac{1}{2}\lambda\phi^{2}}{24\beta^{2}}\right]\qquad(\lambda>0)$ (44)
@@
Following e.g. [27], we now rewrite 44 as
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}\left(1-\frac{T^{2}}{T^{2}_{c}}\right)+\frac{\lambda\phi^{4}}{4!}$ (45)
@@
(a) Symmetry breaking with the effective potential of 45. Here $T_{c}\sim 14.77{\rm K}$ and $T\sim 1{\rm K}$.
@@
Figure 9: Symmetry breaking and restoration of the thermal effective potential of 45 in flat spacetime with $m^{2}{<}0$.
@@
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}\left(1-\frac{T^{2}}{T^{2}_{c}}+\frac{\xi R}{m^{2}}\right)+\frac{\lambda\phi^{4}}{4!}$ (46)
@@
$T_{1c}=\left(1+\frac{\xi R}{m^{2}}\right)^{\frac{1}{2}}T_{c}$ (47)
@@
$V_{\rm{eff,\beta}}=\frac{\lambda\phi^{4}}{4!}+\frac{m^{2}\phi^{2}}{2}\left[1-\frac{T^{2}}{T^{2}_{c}}+\frac{\xi R}{m^{2}}-\frac{3\lambda}{64\pi^{2}}-0.732\times\frac{\left(\frac{1}{6}-\xi\right)R\lambda}{16\pi^{2}m^{2}}\right]$ (48)
@@
$T_{1c}=\left[1+\frac{\xi R}{m^{2}}-\frac{3\lambda}{64\pi^{2}}-0.732\times\frac{\left(\frac{1}{6}-\xi\right)R\lambda}{16\pi^{2}m^{2}}\right]^{\frac{1}{2}}T_{c}$ (49)
@@
Figure 10: Variation of the effective critical temperature with respect to $\lambda$, 49 in the de Sitter spacetime. We have taken $\xi\sim 0.125$ and $\Lambda\sim 10^{-52}{\rm m}^{-2}$.
@@
Since 35 is essentially a high temperature expansion, and we are not setting the mass scale to be equal to $\beta^{-1}$, we will ignore the zero temperature part in 35. In particular, we have explicitly verified that even if we keep them, they bring in no qualitative changes to the symmetry restoration phenomenon at high temperature. Also, in the finite temperature part of the effective action, we have not included the odd power terms of $m_{1}$, as they originate from the zero mode fluctuations (i.e. the $n=0$ term in 28). We refer our reader to [7, 67] and references therein in this context. Nevertheless, we also have explicitly verified that for $\bar{\eta}=0$, inclusion of such terms do not alter any qualitative feature of the symmetry restoration phenomenon.
Let us accordingly begin with the thermal effective potential with zero background spacetime curvature found from 35,
@@
which we have plotted in 11. Comparing this with 3 for zero temperature, we see no qualitative changes. This is expected, as if there is no symmetry breaking at zero temperature, one usually does not expect the same at finite temperature.
@@
Figure 11: Plot of the effective potential at finite temperature$(\beta=10^{-5},\eta=0.447,\lambda=0.1)$ with vanishing background curvature, 50. See main text for discussion.
For the Schwarzschild spacetime, we have from 35
@@
$\left.+0.00001\times(m\beta)^{6}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right]$ (51)
@@
$\frac{V_{\rm{eff},\beta}}{m^{4}}=\frac{\bar{\phi}^{2}}{2}+\frac{\bar{\eta}\bar{\phi}^{3}}{3!}+\frac{\lambda\bar{\phi}^{4}}{4!}+2\xi\Lambda\bar{\phi}^{2}+\left[\frac{\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{24m^{2}\beta^{2}}-0.001\times\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\ln\left((m\beta)^{2}\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right)\right.$
@@
$\left.+0.00004\times(m\beta)^{4}\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)+0.00001\times(m\beta)^{6}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right]$
The variation of the above effective potential can be seen in 13. We have taken $\Lambda\sim 10^{-52}{\rm m}^{-2}$ and $\xi=0.125$. Comparing this with the zero temperature result of 5, we see symmetry restoration. This must be a second order phase transition.
@@
Figure 12: Variation of the effective potential for one loop at finite temperature for the Schwarzschild spacetime, 51.
@@
Figure 13: Plot of the finite temperature effective potential in the de Sitter background ($\beta=10^{-5},\eta=0.447,\lambda=0.1,\Lambda\sim 10^{-52}{\rm m}^{-2}$ and $\xi=0.125$). Comparing this with the zero temperature result of 5, we see symmetry restoration effect at finite temperature and hence indication of a second order phase transition.

### v2

@@ 5 Spontaneous symmetry breaking and phase transition at zero temperature
The chief motivation of this section is to see the effect of the spacetime curvature in the phase transition or spontaneous symmetry breaking phenomenon, at zero temperature. As we have already mentioned, this topic is old. In flat spacetime, one of the earliest thermal field theory calculations in this context can be seen in the seminal work [9]. At zero temperature and with spacetime curvature some such computations can be seen in e.g. [6, 39, 44] and references therein. In particular, [39, 44] use weak or linear in curvature approximation and renormalisation group techniques. However, as we have also mentioned earlier, such results cannot be applicable to Ricci flat spacetimes.
@@
Let us now consider the zero temperature effective potential in the absence of curvature, 21 ($R=0=f_{1}$), in the dimensionless form
@@
$\left.+\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right].$ (41)
Putting in now some customary numerical values, $m\sim 0.0005{\rm GeV}$ and $\mu\sim 3\times 10^{-25}{\rm GeV}$, which may be appropriate for the large scale physics we are interested in, e.g. [87] and references therein, we have plotted the above zero temperature flat space effective potential in 3. We note that since the scalar we have taken has non-vanishing rest mass, there can be no spontaneous symmetry breaking for the quartic self interaction. However, in the presence of spacetime curvature, the scenario can change, as we wish to show below.
@@
Figure 3: The one loop effective potential, 41, with vanishing temperature in flat spacetime. Note that for $\bar{\eta}=0$, no spontaneous symmetry breaking is indicated, as we have taken the rest mass of the scalar to be non-vanishing. Note that by our renormalisation scheme, the effective potential does not contain any field independent finite terms.
@@
in 21, so that we have around $r\sim 2GM$
@@
$\left.+\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right]+\frac{0.0016}{m^{4}G^{4}M^{4}}\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right).$ (42)
Note that the effective potential diverges as $M\to 0$. This corresponds to the fact that the value of $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ increases on the horizon with decreasing mass of the black hole. On the other hand, the result approaches the flat spacetime limit, 41, for large value of the black hole mass. Taking now for example, $GM\sim 10^{2}{\rm m}\sim 10^{17}{\rm GeV}^{-1}$, we have plotted 42 in 4. We note in particular for the first of 4, that for increasing cubic coupling, the effective potential develops a deep. This originates from the logarithm contained in the last term on the right hand side of 42, which may contribute a large negative value. Note also that for a given $\lambda$ value, we cannot arbitrarily go on increasing the cubic coupling, owing to the discussion appearing at the end of the preceding Section, 40.
@@
Figure 4: The one loop effective potential, 42, at vanishing temperature in the Schwarzschild spacetime. See main text for discussion.
@@
$+\left(r^{2}+a^{2}+\frac{2a^{2}GMr\sin^{2}\theta}{r^{2}+a^{2}\cos^{2}\theta}\right)\sin^{2}\theta d\phi^{2},$ (43)
@@
$R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}=\frac{48G^{2}M^{2}\left(r^{6}-15a^{2}r^{4}\cos^{2}\theta+15a^{4}r^{2}\cos^{2}\theta-a^{6}\cos^{6}\theta\right)}{\left(r^{2}+a^{2}\cos^{2}\theta\right)^{6}}.$
Comparing this with the Schwarzschild spacetime, we see that there is no qualitative change at all due to spacetime rotation on the equatorial plane, $\theta=\pi/2$. The effect of the spacetime rotation is maximum on the axis ($\theta=0,\pi$). We note that if we take the extremal limit ($a\to GM$), the Kretschmann scalar vanishes as we move towards the horizon ($r\to GM$) along the axis. This shows that the one loop effective potential for an extreme Kerr black hole spacetime in the vicinity of the axial points near the horizon coincides with that of the flat spacetime, 41.
Finally, we note that neither the Schwarzschild nor the Kerr case shows any SSB. This is because, as can be easily checked, that 42 is always a positive definite for $\bar{\eta}=0$, and the effective potential always must pass through the origin, for all values of the masses $m$ and $M$. Second, changing these mass parameters does not create any new extrema. However, the situation will be qualitatively different for the de Sitter spacetime, as follows.
@@
The metric for the cosmological de Sitter spacetime reads,
$ds^{2}=-dt^{2}+e^{2\sqrt{\Lambda}t/\sqrt{3}}\left(dx^{2}+dy^{2}+dz^{2}\right),$
for which we have $R=4\Lambda$, $R^{\mu\nu}R_{\mu\nu}=4\Lambda^{2}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=8\Lambda^{2}/3$ in 21, $\Lambda$ being the positive cosmological constant. We have the one loop effective potential
@@
$+\frac{0.032\Lambda^{2}}{m^{4}}\ln\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\left[\left(\frac{1}{6}-\xi\right)^{2}-0.001\right].$ (44)
The series appearing in the above equation might seem non-convergent for $\Lambda/m^{2}\ll 1$. This is simply an artefact of our scaling of $V_{\rm{eff}}$ by $m^{4}$. It is easy to see that this apparent ambiguity is absent, should we instead make the scaling with respect to $\Lambda^{2}$. However, since we are primarily dealing with general curved spacetimes, we wish to keep this generic scaling with respect to $m^{4}$. We have investigated various features of this dimensionless effective potential in 5, 6 and 7, for some customary numerical values of different parameters. In particular, we note that (with $\bar{\eta}=0$), spontaneous symmetry breaking at one loop is possible in de Sitter even with $\xi,m^{2}{>}0$. Usually, this has no analogue for flat or the Ricci flat spacetimes we have considered earlier, 3(b), 4(b). In fact symmetry breaking for a massive scalar in flat spacetime seems only to be possible if the mass scale $\mu$ is extremely large compared to $m$. In de Sitter, this effect originates from the term appearing in the last line of the right hand side of 44, which, when $\xi$ is $1/6$ or close to it, yields a negative contribution which can dominate the $\bar{\phi}^{2}/2$ and other positive terms for small values of $\bar{\phi}$, leading to the symmetry breaking effective potential. It is easy to see that this effect will be present even with a tiny value of $\Lambda$. However, such curvature driven symmetry breaking may a priori be expected to be significant in the early inflationary universe scenario, where the dark energy density and hence the spacetime curvature was much higher compared to that of today. Note also from the second of 7 that symmetry breaking remains possible for a very wide range of the renormalisation scale, $\mu$.
@@
Figure 5: The one loop effective potential, 44, at vanishing temperature in the de Sitter spacetime. Comparing the latter with 3(b), 4(b), we see that spontaneous symmetry breaking in de Sitter with a massive scalar with $m^{2}{>}0$ is possible. In both the plots appearing above we have taken $\Lambda\sim 30{\rm Gev}^{2}$, and $\xi\sim 0.13$.
Before we proceed further, a couple of clarifying remarks are in order here. First, we wish to emphasise that the Schwinger-DeWitt expansion formalism based upon the local Lorentz invariance and four momentum we are using here, is essentially meant to capture the curvature effects to ultraviolet quantum processes, such as the anomaly [6]. Being ultraviolet, these are short scale or local effects. Thus, the current formalism, being based upon the local geometry, is not suitable to address the late time, deep infrared or super-Hubble phenomenon like the non-perturbative secular effect in de Sitter [88, 89, 90, 91, 92, 93, 94] (also references therein) for a light scalar. One needs to use the exact propagator and the Schwinger-Keldysh closed time path formalism to address such non-equilibrium scenarios. Thus the effect we are looking into could be some initial or transient phenomenon, with time scale much small compared to that of the duration of the inflation. Thus the present formalism can be thought of as somewhat analogous to the standard particle physics set up, where we essentially ignore the global structure of the spacetime. In fact standard particle physics results or techniques including in-out scattering without considering any curvature have been applied to early inflationary cosmology earlier, provided they are very high energy and short ranged in spacetime, e.g. [95, 96] and references therein. Nevertheless, after the symmetry breaking occurs at short scale, it will be interesting to look into its effect on the background field’s classical or stochastic behaviour at late times and at large scale, from the effective action corresponding to 44.
Finally, we note that had we ignored the quadratic curvature terms from the beginning, there would be no such SSB described above. Will the result still be valid, should we include cubic order curvature terms? Putting things together now, we believe this feature, along with the thermal self energy result 28, sufficiently motivates us to look into this problem via some suitable non-perturbative resummation techniques. Any such technique is not only required to resum over the coupling, but also to resum some curvature terms as well, to see its effect on the short scale physics we are probing here. We reserve these problems for a future work. However in this paper, as a bridge between the above one loop result and a non-perturbative analysis, we wish to present some two loop computations for the effective potential. Our motivation is to see whether the above result gets any qualitative modification at higher loop order. Indeed as we will see, there is no qualitative deviation from SSB in de Sitter. Another motivation of us is to check, in the context of finite temperature field theory, that the field renormalisation counterterms are indeed independent of the spacetime curvature.
@@
Figure 6: The one loop effective potential, 44, at vanishing temperature in the de Sitter spacetime for different values of the non-minimal coupling $\xi$. Note that for $\xi\sim 0.167$ or $1/6$, the dip or the hill of the effective potential is maximum. This effect originates from the terms appearing on the last line on the right hand side of 44. See main text for discussion.
@@
Figure 7: The one loop effective potential, 44, with vanishing temperature in the de Sitter spacetime for different values of the cosmological constant $\Lambda$, and the mass scale $\mu$ (in GeV). Increasing the value of $\Lambda$ increases the strength of the symmetry breaking effect. Note also that the symmetry breaking feature of the potential remains valid for a very wide range of the renormalisation scale, $\mu$.
@@ 5.1 Phase transition at finite temperature
To begin with, let us consider the thermal effective potential 36 at zero spacetime curvature with only quartic coupling and at high temperature, e.g. [33],
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}+\frac{\lambda\phi^{4}}{4!}+\left[-\frac{\pi^{2}}{90\beta^{4}}+\frac{m^{2}+\frac{1}{2}\lambda\phi^{2}}{24\beta^{2}}\right]\qquad(\lambda>0),$ (45)
@@
Following e.g. [33], we now rewrite 45 as
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}\left(1-\frac{T^{2}}{T^{2}_{c}}\right)+\frac{\lambda\phi^{4}}{4!},$ (46)
@@
(a) Symmetry breaking with the effective potential of 46. Here $T_{c}\sim 14.77{\rm GeV}$ and $T\sim 1{\rm GeV}$.
@@
Figure 9: Symmetry breaking and restoration of the thermal effective potential of 46 in flat spacetime with $m^{2}{<}0$.
@@
$V_{\rm{eff},\beta}=\frac{m^{2}\phi^{2}}{2}\left(1-\frac{T^{2}}{T^{2}_{c}}+\frac{\xi R}{m^{2}}\right)+\frac{\lambda\phi^{4}}{4!}.$ (47)
@@
$T_{1c}=\left(1+\frac{\xi R}{m^{2}}\right)^{\frac{1}{2}}T_{c},$ (48)
@@
$V_{\rm{eff,\beta}}=\frac{\lambda\phi^{4}}{4!}+\frac{m^{2}\phi^{2}}{2}\left[1-\frac{T^{2}}{T^{2}_{c}}+\frac{\xi R}{m^{2}}-\frac{3\lambda}{64\pi^{2}}-0.732\times\frac{\left(\frac{1}{6}-\xi\right)R\lambda}{16\pi^{2}m^{2}}\right].$ (49)
@@
$T_{1c}=\left[1+\frac{\xi R}{m^{2}}-\frac{3\lambda}{64\pi^{2}}-0.732\times\frac{\left(\frac{1}{6}-\xi\right)R\lambda}{16\pi^{2}m^{2}}\right]^{\frac{1}{2}}T_{c}.$ (50)
@@
Figure 10: Variation of the effective critical temperature with respect to $\lambda$, 50 in the de Sitter spacetime. We have taken $\xi\sim 0.125$ and $\Lambda\sim 30{\rm GeV}^{2}$.
@@
Since 36 is essentially a high temperature expansion, and we are not setting the mass scale to be equal to $\beta^{-1}$, we will ignore the zero temperature part in 36. In particular, we have explicitly verified that even if we keep them, they bring in no qualitative changes to the symmetry restoration phenomenon at high temperature. Also, in the finite temperature part of the effective action, we have not included the odd power terms of $m_{1}$, as they originate from the zero mode fluctuations (i.e. the $n=0$ term in 29). We refer our reader to [9, 74] and references therein in this context. Nevertheless, we also have explicitly verified that for $\bar{\eta}=0$, inclusion of such terms does not alter any qualitative feature of the symmetry restoration phenomenon.
Let us accordingly begin with the thermal effective potential with zero background spacetime curvature found from 36,
@@
$\left.+5.38\times 10^{-10}(m\beta)^{6}\left(5\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+10\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+10\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}+5\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{4}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{5}\right)\right],$
which we have plotted in 11, ignoring the $\phi$-independent finite terms. Comparing this with 3 for zero temperature, we see no qualitative changes. This is expected, as if there is no symmetry breaking at zero temperature, one usually does not expect the same at finite temperature.
@@
Figure 11: Plot of the effective potential at finite temperature $(\beta=10^{-5}\rm{GeV^{-1}},\eta=0.447,\lambda=0.1)$ with vanishing background curvature, 51. We have not considered finite terms independent of $\phi$. See main text for discussion.
For the Schwarzschild spacetime, we have from 36
@@
$\left.+0.00001\times(m\beta)^{6}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right],$ (52)
@@
$\frac{V_{\rm{eff},\beta}}{m^{4}}=\frac{\bar{\phi}^{2}}{2}+\frac{\bar{\eta}\bar{\phi}^{3}}{3!}+\frac{\lambda\bar{\phi}^{4}}{4!}+\frac{2\xi\Lambda\bar{\phi}^{2}}{m^{2}}+\left[\frac{\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}}{24m^{2}\beta^{2}}-0.001\times\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)\ln\left((m\beta)^{2}\left(1+\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)\right)\right.$
@@
$\left.+0.00004\times(m\beta)^{4}\left(2\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}\right)+0.00001\times(m\beta)^{6}\left(3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)+3\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{2}+\left(\frac{1}{2}\lambda\bar{\phi}^{2}+\bar{\eta}\bar{\phi}\right)^{3}\right)\right].$
The variation of the above effective potential can be seen in 13, after ignoring $\phi$-independent finite terms. We have taken $\Lambda\sim 30{\rm GeV}^{2}$ and $\xi=0.125$. Comparing this with the zero temperature result of 5, we see symmetry restoration. This must be a second order phase transition.
@@
Figure 12: Variation of the effective potential for one loop at finite temperature for the Schwarzschild spacetime, 52.
@@
Figure 13: Plot of the finite temperature effective potential in the de Sitter background ($\beta=10^{-5}\rm{GeV^{-1}},\eta=0.447,\lambda=0.1,\Lambda\sim 30{\rm GeV}^{2}$ and $\xi=0.125$). Comparing this with the zero temperature result of 5, we see symmetry restoration effect at finite temperature and hence indication of a second order phase transition.
The above results are essentially valid at high temperatures. With a low temperature effective potential, we may actually see how the broken symmetry is restored at some critical temperature. Following [9, 33], we keep only the ${\cal O}(\beta^{-2})$ term in this case, so that for de Sitter, we obtain a low temperature version of $V_{{\rm eff},\beta}$,
Figure 14: Variation of the low temperature effective potential, 54, with respect to $\bar{\phi}=\phi/m$, for various values of $\beta$. We have taken $\xi\sim 0.125$ and $\Lambda\sim 30{\rm GeV}^{2}$. Symmetry restoration and the existence of a critical temperature can be explicitly seen here.
$V_{{\rm eff},\beta}\simeq\frac{1}{2}m^{2}\phi^{2}+2\xi\Lambda\phi^{2}+0.032\Lambda^{2}\ln\left(1+\frac{\lambda{\phi}^{2}}{2m^{2}}\right)\left[\left(\frac{1}{6}-\xi\right)^{2}-0.001\right]+\frac{\lambda\phi^{2}}{48\beta^{2}}.$ (54)
For the sake of calculational simplicity, we further take $\phi/m\ll 1$. Then similar steps which led to 50 now yields the critical temperature,
$T_{c}=4.9\sqrt{\frac{\Lambda}{\lambda}}\left[-\frac{m^{2}}{\Lambda}-4\xi+\frac{0.032\lambda\Lambda}{m^{2}}\left(0.001-\left(\frac{1}{6}-\xi\right)^{2}\right)\right]^{1/2},$ (55)
depicted in 14, which is qualitatively similar to that of the $m^{2}<0$ case, discussed in 5.1. It is also clear that a more general analysis requires writing the thermal propagators at general temperatures.